From GlueXWiki
Jump to: navigation, search

May 31, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Production status
    • Fourth package assembly
    • Spare wire planes
  2. Engineering update (Bill)
  3. Electronics update (Chris)
  4. Test set-up in 126 FDC E-log (Beni)
  5. Other


Participants: Bill, Dave, Chris, Nick, Mark, Vlad, Beni, and Lubomir.


- Dave: wire plane #27 remains to be deadened, after tension measurements on #28 will be removed from the stringing table and will string the very last wire frame #8. Chris will put components on #28 next week. Cathode #53 (type-2) had a problem with gluing the ground plane: the Mylar pilled out from the frame. This is the first frame that was converted from type-1 to type-2 by filling the groove with epoxy, so most likely the surface that was epoxied and sanded was too smooth. Casey roughened it and a new foil was tensioned on it today; tomorrow we will know the result.

- Third package was removed from the assembly fixture and put aside on a table. For that all the metal pins had to be removed, then the package was turned upside down to insert the new pins and hubs. We connected the package to the gas to check the tightness: it was as good as it before this operation.

- Today before the meeting we installed the first cell of the fourth package, and after the meeting we installed the second one. The plan is to put also the third cell tomorrow and close the package for flushing and testing. If everything goes smoothly we expect to be ready with the last three cells and install them in the week of June 11-15.

- We started testing the first production wire frames that will be used as spares; all they had some kind of problem. We put the very first (that was submerged in water long time ago) between Lexan sheets and after one day flushing with gas, applied HV. We were able to reach 1500/-500V (to be compared with few hundred volts we could put before) but at this voltage the current started increasing with the time. We decided to stop with the tests, remove all the HV caps and test it again without caps. In case we still have problems then we have to consider making a new wire frame, but we don't have spare PCBs. Chris will look in the old PRs and also will ask the company for an estimate for 4 or 8 PCB sets (each 6 boards), as Bill suggested just to have more spares. For these PCBs we don't want them to put the HV caps.

Engineering update

- Bill: Caleb (the student who worked with him last summer) is back and will be working on the cooling system including its testing. Bill will modify one of the fixtures at Blue Crab to be used for the package spacers.

Electronics update

- Chris on grounding clips: delay with the PR, finally yesterday the clips were ordered, but the delivery time is 3-4 weeks. The third package will have to wait for that, we can't install anything on it before the grounding. Chris will ask them to expedite the order, at least for smaller quantities. According to Nick in ~4 weeks from now we will have all needed to supply one package with LV. Bill: when the cooling system will be needed? - most likely for the fall because we will not have the fADC125 modules before that.

Test set-up at 126

- Beni: No news, done with the tests of the first package and waiting for the third one. In the mean time has to work on two-crate configuration with CODA-light. A student will be working (with the help of Hovanes) on the controls for the testing set-up.


- Gas system: Beni asked Bill if possible to increase the diameter of the gas supply tubes; the impedance looks too high now. Outside of the detector the tube material will be stainless steel to minimize the outgassing. The exhaust line will go outside of the hall as well, to allow use of flammable gases. Beni will discuss this with Narciso Gomez today.